Michael works for a US Senator. She asks Michael to find the inflation rate over the past year to share with a Senate committee that is looking into the size of Cost of Living Adjustments (COLA's) applied to Social Security benefits. Michael finds the following information:

Last Year GDP Deflator 111.1   Consumer Price Index 253.4

This year GDP Deflator 114.3   CPI 258.5

Based on this information, Michael should report to the Senate Committee that the inflation rate was ____________% over the past year.

Answer
Answer #1

Inflation rate for COLA adjustment is measured using CPI. So,

Inflation rate = % Change in CPI = (258.5 / 253.4) - 1 = 1.0201 - 1 = 0.0201 = 2.01%